# FLIP: Towards Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

### v3

@@ V-G Quality of Feature Representations (RQ6)
In this section, we investigate whether the feature-level alignment improves the feature embedding learning and therefore alleviates the representation degeneration problem. Following previous work [36], we adopt SVD [39] decomposition to project the learned feature embeddings of the pretrained ID-based model into 2D data. In Figure 8, we visualize the feature representations learned by four model variants FLIPw/oMLM&MTM, FLIPw/oMLM, FLIPw/oMTM and FLIP. Note that FLIPw/oMLM&MTM indicates the instance-level contrastive learning (ICL) variant without the MLM and MTM pretraining objectives. We have the following observations:
@@
  • The feature embeddings learned by FLIPw/oMLM&MTM (i.e., ICL) collapse into a narrow cone, suffering from severe representation degeneration problem. In contrast, FLIP obtains feature embeddings with more distributed latent patterns, thus better capturing the feature diversity. This highlights the effectiveness of feature-level alignment in promoting representation learning and mitigating the issue of representation degeneration.
@@
  • Comparing FLIP with $\text{FLIP}_{w/oMLM}$ or $\text{FLIP}_{w/oMTM}$, we find that removing either MLM or MTM objective makes the learned embeddings more indistinguishable, demonstrating the necessity of dual alignments between modalities for learning effective representations.
VI Conclusion
@@
In this paper, we propose FLIP, a model-agnostic framework that achieves the fine-grained alignment between ID-based models and PLMs. We view tabular data and transformed textual data as dual modalities, and design a novel modality alignment pretraining task. Specifically, the joint reconstruction for masked language/tabular modeling (with specially designed masking strategies) and cross-modal contrastive learning are employed to accomplish feature-level and instance-level alignments, respectively. Furthermore, we propose to jointly finetune the ID-based model and PLM to achieve superior performance by combining the advantages of both models. Extensive experiments show that FLIP outperforms state-of-the-art baselines on three real-world datasets, and is highly compatible with various ID-based models and PLMs.

### v4

@@ 4.6.2. Visualization
we investigate the impact of the fine-grained alignment on the feature ID embedding learning. Following previous work (Qiu et al., 2022), we adopt SVD (Koren, 2008) decomposition to project the learned ID embeddings from the ID-based model into 2D data. In Figure 6, we visualize the ID embeddings learned by four variants FLIPw/o ${}_{\text{w/o }MLM\&MTM}$, FLIPw/o ${}_{\text{w/o }MLM}$, FLIPw/o ${}_{\text{w/o }MTM}$ and FLIP. Note that FLIPw/o ${}_{\text{w/o }MLM\&MTM}$ indicates the instance-level contrastive learning (ICL) variant without the MLM and MTM objectives. We have the following observations:
@@
  • The feature ID embeddings learned by FLIPw/o ${}_{\text{w/o }MLM\&MTM}$ (i.e., ICL) collapse into a narrow cone, suffering from severe representation degeneration problem. In contrast, FLIP obtains feature ID embeddings with more distributed latent patterns, thus better capturing the feature diversity. This highlights the effectiveness of fine-grained alignment in promoting representation learning and mitigating the representation degeneration.
@@
  • Comparing FLIP with FLIPw/o ${}_{\text{w/o }MLM}$ or FLIPw/o ${}_{\text{w/o }MTM}$, we find that removing either MLM or MTM objective makes the learned embeddings more indistinguishable, demonstrating the necessity of dual alignments between modalities for learning effective representations.
5. Related Work
5.1. ID-based Models for CTR Prediction
CTR prediction serves as a core function module in personalized online services, including online advertising, recommender systems, etc (Zhang et al., 2017). ID-based models follow the common design paradigm: embedding layer, feature interaction (FI) layer, and prediction layer. These models take as input one-hot features of tabular data modality, and employ various interaction functions to capture collaborative signals among features. Due to the significance of FI in CTR prediction, numerous studies focus on designing novel structures for the FI layer to capture more informative and complex feature interactions. Wide&Deep (Cheng et al., 2016) combines a wide network and a deep network to achieve the advantages of both. DCN (Wang et al., 2017) and DCNv2 (Wang et al., 2021) improve Wide&Deep by replacing the wide part with a cross network to learn explicit high-order feature interactions. DeepFM (Guo et al., 2017) combines DNN and FM, and xDeepFM (Lian et al., 2018) extends DeepFM by using a compressed interaction network (CIN) to capture feature interactions in a vector-wise way. Furthermore, explicit or implicit interaction operators are designed to improve performance, such as the productive operator (Qu et al., 2016), the logarithmic operator (Cheng et al., 2020) and the attention operator (Xiao et al., 2017; Song et al., 2019).
5.2. PLMs for CTR Prediction
Pretrained Language Models (PLMs) have demonstrated exceptional success in a wide range of tasks, owing to their extensive knowledge and strong reasoning abilities (Radford et al., 2018, 2019; Brown et al., 2020; Lin et al., 2024b). Inspired by these achievements, the application of PLMs for recommender systems has received more attention (Lin et al., 2023a; Liu et al., 2023; Li et al., 2023c; Zhu et al., 2023; Chen et al., 2023; Fan et al., 2023; Wu et al., 2023; Chen, 2023; Zhu et al., 2023; Wang et al., 2024; Du et al., 2024). Different from one-hot encoding in ID-based models, this line of research needs to convert the raw data into textual modality, thus retaining the original semantic information of features.
Recent efforts to adapt PLMs for CTR prediction have yielded several breakthroughs (Cui et al., 2022; Li et al., 2023b; Bao et al., 2023; Zhang et al., 2023a; Yuan et al., 2023; Wang and Lim, 2023; Ma et al., 2023; Hua et al., 2023; Wei et al., 2023). For instance, CTR-BERT (Muhamed et al., 2021) leverages a two-tower structure with a user BERT and item BERT for final CTR prediction. PTab (Liu et al., 2022) pretrains a BERT model by masked language modeling, and then finetunes it on downstream tasks. P5 (Geng et al., 2022) uniformly converts different recommendation tasks into text generation tasks with T5 backbone (Raffel et al., 2020). However, PLMs encounter challenges in capturing field-wise collaborative signals and discerning features with subtle textual differences. Recent works (Yuan et al., 2023; Li et al., 2023a; Ren et al., 2023) have tried to address this problem by adding text features into the ID-based model or integrating ID information into the PLM. However, they either overlook cross-modal interactions or depend solely on coarse-grained ICL, which is insufficient to capture fine-grained feature-level modality interactions.
Therefore, we propose to conduct fine-grained feature-level alignment between ID-based models and PLMs via the jointly masked tabular/language modeling, which learns fine-grained interactions between tabular IDs and word tokens by the way of mask-and-predict. In the finetuning stage, we also propose to jointly tune both models, and thus leverage the benefits of both textual and tabular modalities to achieve superior CTR prediction performance.
6. Conclusion
@@
In this paper, we propose FLIP, a model-agnostic framework that achieves the fine-grained alignment between ID-based models and PLMs. We view tabular data and transformed textual data as dual modalities, and design a novel fine-grained modality alignment pretraining task. Specifically, the joint reconstruction for masked language/tabular modeling (with specially designed masking strategies) and cross-modal contrastive learning are employed to accomplish feature-level and instance-level alignments, respectively. Furthermore, we propose to jointly finetune the ID-based model and PLM to achieve superior performance by adaptively combining the outputs of both models. Extensive experiments show that FLIP outperforms state-of-the-art baselines on three real-world datasets, and is highly compatible with various ID-based models and PLMs.
Acknowledgements.
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62177033, 62076161) and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102). The work is also sponsored by Huawei Innovation Research Program. We thank MindSpore (min, 2020) for the partial support of this work.